Consider the following balanced reaction

2PbS + 3O2—> 2PbO + 2SO2

rank the following combinations of reactants based on the maximum amount of products that can be produced from the reaction

24.18g lead sulphide and 3.38g oxygen gas

18.97g lead sulphide and 4.02g oxygen gas

22.17g lead sulphide and 3.75g oxygen gas


1
Expert's answer
2020-10-06T09:34:47-0400

Solution:

Number of moles of a substance in a sample is the mass in g divided by the molar mass, which gives the amount in moles.

n = m/M

molar mass of PbS 239.2 gram/mol

molar mass of O2 32.0 gram/mol

According to the reaction equation for 2 mol of PbS (nr PbS) are needs 3 mol of O2 (nr O2) and are forms 2 mol PbO (nr PbO) and 2 mol SO2 (nr SO2).

A.

n PbS = 24.18 g PbS/239.2 g/mol PbS = 0.101 mol PbS

n O2 = 3.38 g O2/32 g/mol O2 = 0.106 mol O2

for 0.106 mol O2 are needs n PbS = 0.106 mol O2 × 2 mol PbS / 3 mol O2 = 0.071 mol PbS.

the limited reactant – O2.

And than, amount of products can be produced

n PbO = n O2 × nr PbO / nr O2

n SO2 = n O2 × nr SO2 / nr O2

n PbO = 0.106 mol O2 × 2 mol PbO / 3 mol O2 = 0.106 mol PbO

n SO2 = 0.106 mol O2 × 2 mol SO2 / 3 mol O2 = 0.106 mol SO2

B.

n PbS = 18.97g PbS/239.2g/mol PbS = 0.079 mol PbS

n O2 = 4.02g O2/32g/mol O2 = 0.126 mol

for 0.079 mol PbS are needs n O2 = 0.079 mol PbS × 3 mol O2 / 2 mol PbS = 0.119 mol O2.

the limited reactant – PbS.

And than, amount of products can be produced

n PbO = n PbS × nr PbO / nr PbS

n SO2 = n PbS × nr SO2 / nr PbS

n PbO = 0.079 mol PbS × 2 mol PbO / 2 mol PbS = 0.079 mol PbO

n SO2 = 0.079 mol PbS × 2 mol SO2 / 3 mol PbS = 0.079 mol SO2

C.

n PbS = 22.17 g PbS/239.2 g/mol PbS = 0.093 mol PbS

n O2 = 3.75 g O2/32 g/mol O2 = 0.117 mol O2

for 0.117 mol O2 are needs n PbS = 0.117 mol O2 × 2 mol PbS / 3 mol O2 = 0.078 mol PbS.

the limited reactant – O2.

And than, amount of products can be produced

n PbO = n O2 × nr PbO / nr O2

n SO2 = n O2 × nr SO2 / nr O2

n PbO = 0.117 mol O2 × 2 mol PbO / 3 mol O2 = 0.078 mol PbO

n SO2 = 0.117 mol O2 × 2 mol SO2 / 3 mol O2 = 0.078 mol SO2


Answer: order of the combinations of reactants: 24.18g lead sulphide and 3.38g oxygen gas; 18.97g lead sulphide and 4.02g oxygen gas; 22.17g lead sulphide and 3.75g oxygen gas.